Second angle of a triangle is 3 times as large as the first angle. The third angle is 4 degrees less than the sum of the 2 angles. Find the measurement of the first angle
Answer by mananth(16946)

Let the first abgle be x
the second angle = 3x
the third angle = x+3x-4
..
x+3x+x+3x-4=180
8x-4=180
8x=184
x=23 deg
..
CHECK
23 + 69 + 23+69-4 =180